OpenStudy (anonymous):

Can someone help me understand this: If f(x) = 2x + 4, find f(3). 10 6 9 4

OpenStudy (anonymous):

The only thing you have to do here, is to replace x by 3, then you get f(3)=2*3+4=6+4=10 :) You're welcome

OpenStudy (anonymous):

10 is your answer!

OpenStudy (anonymous):

Thank you both ! :) I have a few of these that are similar and I was confused. @DrakeNetherlands @TaylorSimonelli

OpenStudy (anonymous):

welcome :) @NANAGX14